OpenStudy (anonymous):

x^3=30 how to put this in calculator properly

OpenStudy (anonymous):

\[\sqrt[3]{30}\]

OpenStudy (anonymous):

that is not coming out correct

OpenStudy (anonymous):

hmmmm, x^3 = 27 ---> x = 3 \[\sqrt[3]{27} = 3\]

OpenStudy (anonymous):

right i am using log functions but i have forgotten how to put them in on the calculator

OpenStudy (anonymous):

thats is the way to answer ur question

OpenStudy (anonymous):

ohhh x^3 = 30 3 ln x = ln 30 ln x = (ln 30)/3

OpenStudy (anonymous):

ln x = 1.133732461 x = e ^1.133732461

OpenStudy (anonymous):

when i put that back in it doesnt solve out to 30 i have logx 30 =3 that i have changed into x^3=30

OpenStudy (anonymous):

nevermind i did not pay attention cool deal thank you very much for your help

OpenStudy (anonymous):

i checked it, x = e ^1.133732461 = 3.107232507 (3.107232507)^3 = 30.000004 = 30 if u need in log way, logx 30 = 3 , in calculator log (30)/log(3), the answer is 29,67304741 round it up 30

OpenStudy (anonymous):

thanks